OpenStudy (anonymous):

The perimeter of a rectangle is six times the width. If the length were increased by 4 inches and the width by 7 inches, the perimeter would be 160 inches. What is the width?

OpenStudy (anonymous):

perimeter is \[P=2l+2w\] and you know this is 6w so put \[6w=2l+2w\] \[4w=2w\] \[2w=l\]

OpenStudy (anonymous):

then put \[2(l+4)+2(w+7)=160\] \[2l+8+2w+14=160\] \[2w+2l+22=160\] \[2w+2l=138\] \[w+l=69\] and since \[l=2w\] this says \[w+2w=69\] \[3w=69\] \[w=13\]

OpenStudy (anonymous):

Back to the top. How did you get 4W=2W?

OpenStudy (anonymous):

Solve the following for L and w:\[\{2L+2w=6w,2(L+4)+2(w+7)=160\}\]\[\{L=46,w=23\}\]First Equation: 2L+2w=6w 2*46 + 2*23 = 6*23 ?, 92 + 46 = 138 ?, 138 = 138 yes Second Equation: 2 (L + 4) + 2 (w + 7) = 160 2*(46 + 4) + 2*(23 + 7) = 160 ?, 2*(50) + 2*(30) = 160 ?, 160 = 160 yes